OpenStudy (anonymous):

If f(x) = (x-1)^3/2 + e^x-2/2, then f^1(2)

myininaya (myininaya):

what is the meaning of ^1? Is that the fist derivative?

OpenStudy (anonymous):

\[f \prime (2)\]

myininaya (myininaya):

Ok, so it is to mean the first derivative. Thanks. So which part are you having trouble finding the derivative of? The first term or the second term?

myininaya (myininaya):

for the first term you can find the derivative of it by using a combination of the chain rule and product rule: \[[(f(x))^n ]'=n[f(x)]^{n-1} \cdot f'(x)\] for the second term you can find the derivative of it using the derivative of e^x is e^x and the chain rule and the constant rule: \[[\frac{e^{f(x)}}{c}]'=\frac{f'(x)e^{f(x)}}{c}\]

myininaya (myininaya):

constant multiple rule*

myininaya (myininaya):

And correct me if I'm wrong but I'm reading it as: \[f(x)=(x-1)^\frac{3}{2}+\frac{e^{x-2}}{2} \]

OpenStudy (anonymous):

yes that is correct